OpenStudy (anonymous):

What is the surface area of a cone with a radius of 9 centimeters and a slant height of 6 centimeters? Use for 3.14 for pi .

OpenStudy (anonymous):

SA = πr^2 + πrl

OpenStudy (anonymous):

?

OpenStudy (anonymous):

That's the formula for surface area. l = slant height

OpenStudy (anonymous):

SA = π(9)^2 + π(r)(6)

OpenStudy (anonymous):

How would I do it from there?

OpenStudy (anonymous):

Multiply it out? SA = π(9)^2 + π(9)(6) SA = 81π + 54π SA = ?